NASA communicates with the Space Shuttle and International Space Station using Ku-band microwave radio. Suppose NASA transmits a microwave signal to the Space Shuttle using radio waves with a frequency of 18.GHz. Calculate the wavelength of these radio waves. Round your answer to 2 significant digits.
Answer #1

Given:

f = 18 GHz = 1.8*10^10 s-1

use:

lambda = c/f

=(3.0*10^8 m/s)/(1.8*10^10 s-1)

= 1.67*10^-2 m

= 16.7 mm

Answer: 17 mm